OpenStudy (anonymous):

Arithmetic question: 7/20

OpenStudy (anonymous):

\[\bf \frac{ 7 }{ 20 }=\frac{ 7(5) }{ 20(5) }=\frac{ 35 }{ 100 }=0.35\]

OpenStudy (anonymous):

@genius12 how did you got 5

OpenStudy (bahrom7893):

he multiplied by \[1=\frac{5}{5}\]

OpenStudy (loser66):

because you need 100 at denominator to get the answer. So, 5 is the right number to get that and you cannot multiply 5 just to denominator, but both them: numerator and denominator. That's why genius12 do that step